# Fast rotations in galaxies at cosmic noon indicate central concentration of stars, dark matter or massive black holes

### v1

@@ 5.1.4 Alternative to the assumed dark matter distributions
The dynamical models presented so far have been constructed under the assumption that dark matter follows an NFW density profile. However, alternative halo models have been proposed, and although we do not explore them quantitatively here, they are worth briefly discussing. One possibility is that the dark matter halo profile has been modified by adiabatic contraction due to the condensation of baryons in the central regions of galaxies (e.g., Blumenthal_1986; gnedin_2004; Hussein_2025). As gas cools and settles into the galactic disc and bulge, it deepens the gravitational potential well, causing the dark matter to contract and become more centrally concentrated. This process can steepen the inner slope of the halo relative to the NFW profile (li2022).
Another possibility arises from dark matter models that are alternatives to CDM. Particle physicists have proposed a variety of dark matter candidates in the dark sector that can lead to departures from the NFW shape (Tulin_2018). In particular, self-interacting dark matter (SIDM) models predict a wide diversity of halo structures, ranging from cored to even more cuspy profiles than NFW (e.g., despali19; correa25). Given the limited number statistics of our sample, we defer a detailed investigation of these alternative models to future work, when improved constraints on the baryonic distributions will also be available (e.g., from the ALMA–CONDOR sample111https://almascience.eso.org/alma-data/lp).
@@ 5.2 Dark matter fractions
In recent years, several studies based on warm gas observations have estimated the dark matter fraction within the effective radius, , comparing it with simulations and with other galaxy properties, such as stellar surface density and star formation rate (genzel17; bouche22; puglisi23; degraaff24). The relatively low dark matter fractions , typically found in massive galaxies, have often been interpreted as evidence for the presence of a central dark matter core (genzel20; nestor23; Danhaive_2025b). These claims have motivated theoretical works to explore various mechanisms for the formation of such cores at , including the impact of dry mergers combined with AGN-driven feedback (Dekel_2021), or the infall of baryonic clumps (Ogiya_2022). However, these observational studies rely on several assumptions, such as warm gas accurately tracing circular motions and on gas fractions and distributions inferred from scaling relations (see Section 1). Moreover, it is worth stressing that nearby massive galaxies typically have low dark matter fractions but cuspy halos (e.g. mancerapina22b).
@@
We compare our values with the KURVS-CDFS sample from puglisi23, the IRAM-NOEMA/RC41 sample from genzel20, and the two galaxies from lelli23. For reference, we also include local spiral galaxies from the SPARC survey, which contains 175 nearby disc galaxies with high-quality rotation curves traced primarily by H I (and H in the inner regions of massive systems) and 3.6 m photometry (lelli16). Finally, we include the trend for simulated galaxies in EAGLE at (crain15). We find that the ALPAKA galaxies, all with dark matter fractions of about 0.2 within the disc effective radius, are broadly consistent with values reported both for local galaxies and, importantly, for systems at –. However, our interpretation differs substantially from previous studies. Based on our fiducial models of the rotation curve decomposition, these galaxies are well described by standard NFW halos. In fact, if anything, they may suggest dark matter halos that are cuspier than the NFW profile (Section 5.1.4).
@@ 6 Conclusions
In this work, we followed up on three cosmic noon ALPAKA galaxies (ID1, ID3 and ID13) with both highly resolved CO/[CI] emission from ALMA, covering the cold gas, and rest-frame near infrared JWST data, covering the bulk of the stellar populations. We used the rotation curves obtained with 3DBAROLO in rizzo23, corrected for pressure support and modelled the mass contribution of the baryonic and dark matter components in the galaxies. We assumed that the baryons are distributed in an exponential gas disc, an exponential stellar disc and a spherical bulge, and that the dark matter is distributed in an NFW halo.
@@ Acknowledgements
We thank Enrico Di Teodoro for valuable discussions. FRO and FF acknowledges support from the Dutch Research Council (NWO) through the Klein-1 Grant code OCEN2.KLEIN.088. FR acknowledges support from the Dutch Research Council (NWO) through the Veni grant VI.Veni.222.146. This work has received funding from the European Research Council (ERC) under the Horizon Europe research and innovation programme (Acronym: FLOWS, Grant number: 101096087). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.01659.S; ADS/JAO.ALMA#2017.1.00471.S; ADS/JAO.ALMA#2017.1.01674.S; ADS/JAO.ALMA#2015.1.00862.S; ADS/JAO.ALMA#2017.1.01228.S; ADS/JAO.ALMA#2018.1.00974.S; ADS/JAO.ALMA#2017.1.00413.S; ADS/JAO.ALMA#2019.1.01362.S; ADS/JAO.ALMA#2017.1.01045.S; ADS/JAO.ALMA#2013.1.00059.S; ADS/JAO.ALMA#2018.1.00543.S; ADS/JAO.ALMA#2015.1.00723.S; ADS/JAO.ALMA#2018.1.01146.S; ADS/JAO.ALMA#2016.1.01155.S; ADS/JAO.ALMA#2017.1.01677.S; ADS/JAO.ALMA#2018.1.01306.S; ADS/JAO.ALMA#2017.1.00908.S; 2023.1.00714.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We acknowledge usage of the Python programming language (vanrossum09), Astropy (astropycollab22), Matplotlib (hunter07), NumPy (vanderwalt11), and SciPy (virtanen20).
@@
The JWST data underlying this article are publicly available and were accessed from the DAWN JWST Archive. The ALMA data are publicly available and can be retrieved through the ALMA Archive website.

### v2

@@ 5.1.4 Alternative to the assumed dark matter distributions
The dynamical models presented so far have been constructed under the assumption that dark matter follows an NFW density profile. However, alternative halo models have been proposed, and although we do not explore them quantitatively here, they are worth briefly discussing. One possibility is that the dark matter halo profile has been modified by adiabatic contraction due to the condensation of baryons in the central regions of galaxies (e.g., Blumenthal et al., 1986; Gnedin et al., 2004; Hussein et al., 2025). As gas cools and settles into the galactic disc and bulge, it deepens the gravitational potential well, causing the dark matter to contract and become more centrally concentrated. This process can steepen the inner slope of the halo relative to the NFW profile (Li et al., 2022).
Another possibility arises from dark matter models that are alternatives to CDM. Particle physicists have proposed a variety of dark matter candidates in the dark sector that can lead to departures from the NFW shape (Tulin and Yu, 2018a). In particular, self-interacting dark matter (SIDM) models predict a wide diversity of halo structures, ranging from cored to even more cuspy profiles than NFW (e.g., Despali et al., 2019; Correa et al., 2025). Given the limited number statistics of our sample, we defer a detailed investigation of these alternative models to future work, when improved constraints on the baryonic distributions will also be available (e.g., from the ALMA–CONDOR sample111https://almascience.eso.org/alma-data/lp).
@@ 5.2 Dark matter fractions
In recent years, several studies based on warm gas observations have estimated the dark matter fraction within the effective radius, , comparing it with simulations and with other galaxy properties, such as stellar surface density and star formation rate (Genzel et al., 2017; Bouché et al., 2022; Puglisi et al., 2023; de Graaff et al., 2024). The relatively low dark matter fractions , typically found in massive galaxies, have often been interpreted as evidence for the presence of a central dark matter core (Genzel et al., 2020; Nestor Shachar et al., 2023; Danhaive et al., 2025a). These claims have motivated theoretical works to explore various mechanisms for the formation of such cores at , including the impact of dry mergers combined with AGN-driven feedback (Dekel et al., 2021), or the infall of baryonic clumps (Ogiya and Nagai, 2022). However, these observational studies rely on several assumptions, such as warm gas accurately tracing circular motions and on gas fractions and distributions inferred from scaling relations (see Section 1). Moreover, it is worth stressing that nearby massive galaxies typically have low dark matter fractions but cuspy halos (e.g. Mancera Piña et al., 2022a).
@@
We compare our values with the KURVS-CDFS sample from Puglisi et al. (2023), the IRAM-NOEMA/RC41 sample from Genzel et al. (2020), and the two galaxies from Lelli et al. (2023). For reference, we also include local spiral galaxies from the SPARC survey, which contains 175 nearby disc galaxies with high-quality rotation curves traced primarily by H I (and H in the inner regions of massive systems) and 3.6 m photometry (Lelli et al., 2016b). Finally, we include the trend for simulated galaxies in EAGLE at (Crain et al., 2015). We find that the ALPAKA galaxies, all with dark matter fractions of about 0.2 within the disc effective radius, are broadly consistent with values reported both for local galaxies and, importantly, for systems at –. However, our interpretation differs substantially from previous studies. Based on our fiducial models of the rotation curve decomposition, these galaxies are well described by standard NFW halos. In fact, if anything, they may suggest dark matter halos that are cuspier than the NFW profile (Section 5.1.4).
@@ 6 Conclusions
In this work, we followed up on three cosmic noon ALPAKA galaxies (ID1, ID3 and ID13) with both highly resolved CO/[CI] emission from ALMA, covering the cold gas, and rest-frame near infrared JWST data, covering the bulk of the stellar populations. We used the rotation curves obtained with 3DBAROLO in Rizzo et al. (2023), corrected for pressure support and modelled the mass contribution of the baryonic and dark matter components in the galaxies. We assumed that the baryons are distributed in an exponential gas disc, an exponential stellar disc and a spherical bulge, and that the dark matter is distributed in an NFW halo.
@@ Acknowledgements
We thank Enrico Di Teodoro for valuable discussions. FRO and FF acknowledges support from the Dutch Research Council (NWO) through the Klein-1 Grant code OCEN2.KLEIN.088. FR acknowledges support from the Dutch Research Council (NWO) through the Veni grant VI.Veni.222.146. This work has received funding from the European Research Council (ERC) under the Horizon Europe research and innovation programme (Acronym: FLOWS, Grant number: 101096087). ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The JWST data underlying this article are publicly available and were accessed from the DAWN JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.
@@
The JWST observations used in this paper are publicly available and are associated with program #1180, #1727, #1837 and can be obtained from https://dx.doi.org/10.17909/6qq5-w568. The ALMA data are publicly available and can be retrieved through the ALMA Archive website. The ALMA data are associated with the following programs: ADS/JAO.ALMA#2017.1.01659.S; ADS/JAO.ALMA#2017.1.00471.S; ADS/JAO.ALMA#2016.1.01426.S; ADS/JAO.ALMA#2017.1.00413.S.